# Code-Driven Law NO, Normware SI!

### v1

@@ Abstract
With the digitalization of society, the interest, the debates and the research efforts concerning “code”, “law”, “artificial intelligence”, and their various relationships, have been widely increasing. Yet, most arguments primarily focus on contemporary computational methods and artifacts (inferential models constructed via machine-learning methods, rule-based systems, smart contracts, …), rather than attempting to identify more fundamental mechanisms. Aiming to go beyond this conceptual limitation, this paper introduces and elaborates on “normware” as an explicit additional stance — complementary to software and hardware — for the interpretation and the design of artificial devices. By means of a few examples, we argue that normware-centred views provide a more adequate abstraction to study and design interactions between computational systems and human institutions, and may help with the design and development of technical interventions within wider socio-technical views.
@@ 1 Introduction
The concept of code-driven law, ie. of “legal norms or policies that have been articulated in computer code” by some actors with normative competence, has been convincingly elaborated by Hildebrandt in recent work [1]. Its introduction has the merit to refocus the discussion on the role of artificial devices in the legal activity, rather than on ontological positions expressed under code-is-law or law-is-code banners, which are present, with various interpretations and changing fortunes, in the literature and practice of contemporary regulatory technologies, and technology-oriented legal scholarship (see the overview in [2]).
@@
According to Hildebrandt, code-driven law should be distinguished both from data-driven law, ie. computational decision-making derived from statistical or other inductive methods (including predictions used in support for human decision-making), and from text-driven law, ie. the legal activity performed by humans by means of sources of norms such as statutory and case law. A crucial difference between these forms of “law” is that the linguistic artifacts used in text-driven law are characterized by open-textured concepts (eg. reasonableness, bona fide, foreseeability, etc.) and multi-interpretability, and therefore carry inherently a potential of dispute that data-driven or code-driven law do not have (for statistical, or for logical closures). For this reason, code-driven law could lead at best to computational legalism [3], promoting legal certainty, but not justice (proportionality, distribution, etc.), nor instrumentality (ie. taking into account policy goals set by the legislature). In Hildebrandt’s words: “What code-driven law does is to fold enactment, interpretation and application into one stroke, collapsing the distance between legislator, executive and court.”
@@
The present paper aims to respond to Hildebrandt’s analysis, by reframing some of the identified issues — which are definitively observable in contemporary technologies — in terms of technological gaps. What if we allow enactment, interpretation and application to be unfolded by design? Motivated by this question we will form a more nuanced view of the problem, highlighting spaces of computational research and practice that are (so far) for the most neglected, or partially addressed in different tracks, without enjoying the benefits of a unifying vision. In particular, the paper presents several arguments supporting the introduction of “normware” as an explicit third level — besides hardware and software — for the conceptualization and the design of artificial devices. [^1] The term “normware” was inspired by works on computational normative systems, traditionally focusing on legal knowledge artifacts. With this meaning, it was introduced in the workshop “Normware: Modeling Norms and Automated Norm Application”, held at the Lorentz Center in Leiden (2016). Yet, we soon acknowledged that the traditional symbolic approach was insufficient to deal with structural problems raised e.g. by explainable AI and trustworthy AI [4], and more in general with efforts concerning responsible computing, as for instance those addressed by the Responsible Internet proposal [5]. This paper widely extends and complements these lines of discussions, constructing an overarching theoretical framework which is only sketched in previous works.
The contribution of this work, although concerning a technical domain, is primarily conceptual. Its intended audience are both computer scientists working at the interface between computational and human systems, as well as legal scholars working on technological dimensions of society.
@@
Readers, in particular those with a humanistic background, may find connections, echoes or at least analogies of parts of this exposition to seminal conceptual frameworks of authors as eg. Deleuze’s assemblage [6], Luhmann’s theory of social systems [7], Latour’s Actor-Network theory [8]. This is not surprising: although our main goal here is identifying and mapping engineering dimensions relevant to the regulation of computational systems, the overall effort can also be seen as formulating aspects of social systems, albeit computational. Yet, the perspective and the assumptions underpinning the present work are rather different from those taken by these authors. We start here from engineering concerns, and therefore ontological considerations — figuring an infrastructure, ie. reusable computational components and their mechanisms of interactions — are much stronger than epistemological ones. In purpose, this paper should be rather connected to contributions as Newell’s knowledge level [9] (1982), which, reorganizing in an accessible conceptual framework various computational approaches based on reified explicit knowledge, paved the way to various achievements of symbolic AI, including the semantic web technology. As the paper will argue for, the normware level goes beyond (or rather encompasses) the knowledge level, as it does not presuppose a symbolic reification, nor an intentional unity: it is meant to capture the more diverse forms of normativity expressed through computational systems, either reified as artifacts, or realized as processes.
@@
The paper proceeds as follows. Section 2 starts from revisiting Hildebrandt’s distinction [1] between text-driven, data-driven, and code-driven law. I argue that the differences between the three, in particular in terms of contestability, may be less pronounced than what previously thought. From an engineering perspective, this observation opens up the opportunity to reframe the problem differently, namely by introducing the notion of normware. Section 3 will elaborate on normware as an artifact, taking both internal and external views, and identifying its three main functions: regulating behaviours, qualifications, and expectations. Section 4 will focus instead on normware as a process, filling an existing gap between the traditional “control” stance of hardware and software engineering to the “guidance” stance of policy- and norm-making. Section 5 will reply to possible objections derived from reductionist stances (“normware is just software”), offering examples of application of normware as interpretative stance. Section 6 delineates research tracks whose relevance is further motivated by an explicit distinction (“normware is not software”).
@@ 2.1 Text-guided law?
In recent work [1], Hildebrandt refers to “text-driven law” as the legal activity performed by humans by means of sources of norms expressed in verbal forms. Such an activity could not have been performed without a whole—historically non-computational— technological infrastructure related to textual artifacts (used for printing, distributing, archiving, methods for indexing and retrieving) [10]. However, at a more fundamental level, text-based law inherently relies on language, and natural language functions differently than formal languages used in computational settings. Natural language is open to polysemy, it strongly relies on analogical constructs, and its use (in production and in interpretation) heavily depends on contextualization. In some occasion, text in laws is intentionally left ambiguous, for instance when legislators, also for reasons of political opportunity, prefer not to take a clear commitment towards a certain outcome. In other cases, multi-interpretability is a consequence of generalization. Law is expected to work in a differed fashion, it does not apply only on known, expected cases, but it aims to be relevant also for cases that have not yet appeared. Humans in the loop are therefore inevitable, in particular with cases exhibiting new features or when changes are made to the institutional structure. Multi-interpretability also entails that different humans may frame the same case selecting and interpreting texts in different ways, opening to disputes and also to uncertainty with respect to the outcome of the decision. Part of the procedural mechanisms in court as well as of the jurisprudential debate have a crucial role in promoting legal certainty, but what matters here to us is that the relation between text and legal activity is not a direct one, and, insofar humans are involved, we should rather refer to “text-guided law”.222 Following a similar idea, if data-driven or code-driven inferential mechanisms are part of a support system for providing information or recommendations to human decision-makers on legal matters, it would be more proper to talk of “data-guided law” or “code-guided law”.
@@ 2.2 Uncertainties in data-driven law
Regardless of how data-driven inferential mechanisms are incorporated in the decision-making process, empirical alignment resulting from machine learning is generally performed against a certain training data-set, in principle collected and engineered to provide a sound statistical closure to the target domain of application. The effectiveness of the trained model to achieve the desired outcome depends not only on the dataset, but also on the specific training method. Let us suppose, without loss of generality, that the data-set consists of a set of instances, each one associated with a label, as in a traditional supervised classification task. From a technical point of view, it is well known that good practices in data science and machine learning associate with critical questions, amongst which: Are the instances in the training data set representative of the domain upon which the model is applied? Is the rationale behind the labelling still the same at runtime? Is the chosen training method providing adequate, reproducible, robust results? The third question can be directly associated to technical aspects, but the first two are prototypical examples of context-centred (and not algorithm-centred) inquiries, in particular concerning the assumptions of statistical closure. [^3] See e.g. the RETRO-VIZ tool [11], estimating and explaining the trustworthiness of predictions on the basis of the statistical grounds on which the model has been constructed. Thus, data-driven law can be contested, and indeed best practices entail that it should be scrutinized, even before setting additional requirements as eg. explainability or transparency of the inferential process.
@@ 2.3 Uncertainties in code-driven law
The third form of legal activity identified by Hildebrandt is “code-driven law”, for which provisions from various legal sources (eg. legal norms, regulations, policies, agreements, contracts) have been operationalized via computational means. From a historical point of view, the idea of using machines with normative provisions traditionally dates back to Leibniz (Calculemus!), and took more concrete forms with the advent of legal expert systems and various attempts at formalisation of law (eg. [12, 13]), in part renewed during the semantic web technology heyday, and, in narrower forms, in contemporary RegTech (eg. smart contracts), and cyber infrastructure research (eg. data governance, consent management systems). Most distinctive assumptions of these contributions can be categorized within the domain of legal isomorphism [14], susceptible to concerns on whether, to what extent, and how legal sources can be represented in computational artifacts, and normative reasoning automated by computational processes. At a strictly theoretical level, neglecting interpretative and tractability challenges, the problem becomes a purely logical one. Given a formal language, and the associated derivation mechanisms, soundness and completeness can be formally proven to verify a logical closure guaranteeing valid inferences. Let us then suppose that our system is provided with structural knowledge relative to norms in the form of rules, and contingent knowledge relative to the case in the form of facts; within a formal system we can be certain that we utilize these resources to derive logically valid conclusions. But are they correct in a “practical” sense? Theory and practice of computational legal theory acknowledge several critical questions in this respect: Were all the relevant fact and rules considered? Is the formalization correct? Is the reasoning tool robust? As in the previous section, only the last question concerns proper technical dimensions (eg. tractability, verification, security); the other two concern the selection of relevant contextual elements, and the choice of mapping from natural to formal language.
@@ 2.4 Setting an alternative technological agenda
Taking into account the critical questions identified in the respective fields of practice, we can conclude that neither data-driven nor code-driven law guarantee legal certainty. The premises on which the underlying computational devices are constructed and deployed will always be “weak” to some extent. Besides, software bugs, limitations or hacks (malicious attacks) may occur in computational systems supporting both. Said differently, the computational legalism “ideal” [^4] We refer to Diver [3] for a detailed inquiry on the concept of computational legalism. Related concepts can also be found in early works in AI & Law, cf. mechanical jurisprudence in [15], or the “legalistic child” example discussed in [16]. works only in so far we forget all what may go wrong. The only way by which these computational devices may achieve legal certainty is by removing the possibility of contestation, which, although this may be seen as a feature of the system, is clearly societally unacceptable.
@@
Perhaps, a more sustainable path is to reproduce in the computational realm procedures analogous to the ones introduced in human societies, which have proven value in the course of civilizations to deal with the legal (un)certainty of text-guided law. For instance, in order to satisfy principles of justice, such as proportionality and distribution, we may formalize those as meta-rules. In order to keep instrumentality with respect to contextual policy goals, we may require to make these goals explicit, and have them in the loop of the automated reasoning process. The resulting (naive) computational justice stance is based on the belief that a perfect conceptualization will solve all problems observed before. Unfortunately, there are two serious hindrances for such an approach. First, we cannot be sure of any formalization, ever. The strongest support for this claim is not technical, but comes from the human dimension of the problem, the “referent” (the entity referred to) that needs to be formalized. For instance, the debates on ethical frameworks and human rights make clear that there is no such a referent, or, if it exists, it does so only on some contextual basis. [^5] See e.g. Schlag’s metaphor of law as “hidden ball” [17]. Second, we may never be ready to have an explicit and complete articulation of policy-makers’ purposes. Even imaging that this was possible, by means of perfect introspection and no errors in expression, making clear in detail all what some authority aims to achieve (or does not aim to) would entail to be transparent also on highly contestable issues, which in many occasions would be politically detrimental, and thus not to be expected.
@@
When both computational legalism and (naive) computational justice have been defeated, it seems there is no much hope for an advance in legal technologies. Fortunately, this is not true; new and sounder objectives can be set by starting from opposite assumptions. With respect to computational legalism, we should not suppose that things are correct by default (or eventually will be), but instead accept that things may always go wrong, at all levels. Under this new expectation, we may still accept to use legal technologies, as long as we accept that also that people or other actors ought to be able to appeal algorithmic decisions. Similarly, with respect to computational justice, we should accept that there may always be relevant directives not adequately articulated in the present computational artifacts. This entails that we need means to integrate systematically feedbacks from higher-order courts, from jurisprudence or from other competent normative sources. In order words, courts or other actors ought to be able to quash or overrule algorithmic decision-making processes. Accepting these two principles leads us to technological innovations which have been neglected so far, addressing primarily computational instruments that operationalize appeal and quashing/overruling within the computational realm. Interestingly, within such an infrastructure, we can systematically introduce artificial agents, additionally to humans, to perform critical tasks of testing and auditing.
In short, the core challenge of regulatory technologies is not captured by the requirements expressed in software engineering terms, but by the construction of proper, functioning feedback loops, in the form of intervention points for social actors provided with relevant competence. At this level of abstraction, we are in front of a problem of governance (who has the power to do what, under which circumstances, and for what purpose), a problem which, when computational systems are involved as mediators, has also a technical dimension to be dealt with. This is where the concept of normware fits into.

### v2

@@ Abstract
With the digitalization of society, the interest, the debates and the research efforts concerning “code”, “law”, “artificial intelligence”, and their various relationships, have been widely increasing. Yet, most arguments primarily focus on contemporary computational methods and artifacts (inferential models constructed via machine-learning methods, rule-based systems, smart contracts), rather than attempting to identify more fundamental mechanisms. Aiming to go beyond this conceptual limitation, this paper introduces and elaborates on “normware” as an explicit additional stance — complementary to software and hardware — for the interpretation and the design of artificial devices. By means of a few examples, I will argue that a normware-centred perspective provides a more adequate abstraction to study and design interactions between computational systems and human institutions, and may help with the design and development of technical interventions within wider socio-technical views.
@@ 1 Introduction
The concept of code-driven law, i.e. of “legal norms or policies that have been articulated in computer code” by some actors with normative competence, has been convincingly elaborated by Hildebrandt [1]. Its introduction has the merit to refocus the discussion on the role of artificial devices in the legal activity, rather than on ontological positions expressed under code-is-law or law-is-code banners, which are present, with various interpretations and changing fortunes, in the literature and practice of contemporary regulatory technologies, and technology-oriented legal scholarship (see the overview in [2]).
@@
According to Hildebrandt, code-driven law should be distinguished from data-driven law, i.e. computational decision-making derived from statistical or other inductive methods, and from text-driven law, i.e. the legal activity performed by humans by means of sources of norms such as statutory and case law. A crucial difference between these forms of “law” is that the linguistic artifacts used in text-driven law are characterized by open-textured concepts (e.g. reasonableness, bona fide, foreseeability, etc.) and multi-interpretability, and therefore carry inherently a potential of dispute that data-driven or code-driven law do not have (for statistical, or for logical closures). For this reason, code-driven law could lead at best to computational legalism [3], promoting legal certainty, but not justice (proportionality, distribution, etc.), nor instrumentality (i.e. taking into account policy goals set by the legislature). In Hildebrandt’s words: “What code-driven law does is to fold enactment, interpretation and application into one stroke, collapsing the distance between legislator, executive and court.”
@@
The present paper aims to respond to Hildebrandt’s analysis, by reframing some of the identified issues — which are definitively observable in contemporary technologies — in terms of technological gaps. What if we allow enactment, interpretation and application to be unfolded by design? Motivated by this question we will form a more nuanced view of the problem, highlighting spaces of computational research and practice that are (so far) for the most neglected, or partially addressed in different tracks, without enjoying the benefits of a unifying vision. In particular, the paper presents several arguments supporting the introduction of “normware” as an explicit third level — besides hardware and software — for the conceptualization and the design of artificial devices. [^1] The term “normware” was initially inspired by works on computational normative systems, traditionally focusing on legal knowledge artifacts. With this meaning, it was introduced in the workshop “Normware: Modeling Norms and Automated Norm Application”, held at the Lorentz Center in Leiden (2016). Yet, we soon acknowledged that the traditional symbolic approach was insufficient to deal with structural problems raised e.g. by explainable AI and trustworthy AI [4], and more in general with efforts concerning responsible computing, as for instance those addressed by the Responsible Internet proposal [5]. This paper widely extends and complements these works, constructing an overarching theoretical framework which was previously only sketched.
@@
The contribution of this work, although concerning a technical domain, is primarily conceptual. Its intended audience are both computer scientists working at the interface between computational and human systems, as well as legal scholars working on technological dimensions of society. Readers may find connections, echoes or at least analogies of parts of this exposition to seminal conceptual frameworks of authors as e.g. Deleuze’s assemblage [6], Luhmann’s theory of social systems [7], Latour’s Actor-Network theory [8]. This is not surprising: although our main goal here is identifying and mapping engineering dimensions relevant to the regulation of computational systems, the overall effort can also be seen as formulating aspects of social systems, albeit computational. Yet, the perspective and the assumptions underpinning the present work are rather different from those taken by these authors. We start here from engineering concerns, and therefore ontological considerations — figuring an infrastructure, i.e. reusable computational components and their mechanisms of interactions — are much stronger than epistemological ones. In purpose, this paper should rather be connected to contributions as Newell’s knowledge level [9] (1982), which, reorganizing in an accessible conceptual framework various computational approaches based on reified explicit knowledge, paved the way to various achievements of symbolic AI, including the semantic web technology. As we will elaborate in the paper, the normware level encompasses the knowledge level, because it does not presuppose a symbolic reification, nor an intentional unity: it is meant to capture the more diverse forms of normativity expressed through computational systems, either reified as artifacts, or realized as processes.
@@
The paper proceeds as follows. Section 2 starts from revisiting Hildebrandt’s distinction [1] between text-driven, data-driven, and code-driven law. The differences between the three, in particular in terms of contestability, may be less pronounced than what previously thought. From an engineering perspective, this observation opens up to reframing the problem differently, and we do so by introducing the notion of normware. Section 3 will elaborate on normware as an artifact, taking both internal and external views, and identifying its three main functions: regulating behaviours, qualifications, and expectations. Section 4 will focus instead on normware as a process, filling an existing gap between the traditional “control” stance of hardware and software engineering to the “guidance” stance of policy- and norm-making. Section 5 will reply to possible objections derived from reductionist stances (“normware is just software”), offering examples of application of normware as interpretative stance. Section 6 delineates research tracks whose relevance is further motivated by an explicit distinction (“normware is not software”).
@@ 2.1 Text-guided law?
In recent work [1], Hildebrandt refers to “text-driven law” as the legal activity performed by humans by means of sources of norms expressed in verbal forms. Such an activity could not have been performed without a whole—traditionally non-computational—technological infrastructure related to textual artifacts, used for printing, distributing, archiving, methods for indexing and retrieving [10]. Yet, at a more fundamental level, text-based law inherently relies on language, and natural language functions differently than formal languages used in computational settings. Natural language is open to polysemy, it strongly relies on analogical constructs, and its use (in production and in interpretation) heavily depends on context. In some occasion, text in laws is intentionally left ambiguous, for instance when legislators, for reasons of political opportunity, prefer not to take a clear commitment towards a certain outcome. In other cases, multi-interpretability is a consequence of generalization. Law is expected to work in a differed fashion, it does not apply only on known, expected cases, but it aims to be relevant also for cases that have not yet appeared. Humans in the loop are therefore inevitable, in particular with cases exhibiting new features or when changes are made to the institutional structure. Multi-interpretability also entails that different humans may frame the same case selecting and interpreting texts in different ways, opening to disputes and also to uncertainty with respect to the outcome of the decision. Part of the procedural mechanisms in court as well as of the jurisprudential debate have a crucial role in promoting legal certainty, but what matters here to us is that the relation between text and legal activity is not a direct one, and, insofar humans are involved, we should rather refer to “text-guided law”.222 Following a similar idea, if data-driven or code-driven inferential mechanisms are part of a support system for providing information or recommendations to human decision-makers on legal matters, it would be more proper to talk of “data-guided law” or “code-guided law”.
@@ 2.2 Uncertainties in data-driven law
Regardless of how data-driven inferential mechanisms are incorporated in the decision-making process, empirical alignment resulting from machine learning is generally performed against a certain training data-set, in principle collected and engineered to provide a sound statistical closure to the target domain of application. The effectiveness of the trained model to achieve the desired outcome depends not only on the dataset, but also on the specific training method. Let us suppose, without loss of generality, that the data-set consists of a set of instances, each one associated with a label, as in a traditional supervised classification task. From a technical point of view, it is well known that good practices in data science and machine learning associate with critical questions, amongst which: Are the instances in the training data set representative of the domain upon which the model is applied? Is the rationale behind the labelling still the same at runtime? Is the chosen training method providing adequate, reproducible, robust results? The third question can be directly associated to technical aspects, but the first two are prototypical examples of context-centred (and not algorithm-centred) inquiries, in particular concerning the assumptions of statistical closure. [^3] See e.g. the RETRO-VIZ tool [11], estimating and explaining the trustworthiness of predictions on the basis of the statistical grounds on which the model has been constructed. Thus, data-driven law can be contested, and indeed best practices entail that it should be scrutinized, even before setting additional requirements as e.g. explainability or transparency of the inferential process.
@@ 2.3 Uncertainties in code-driven law
The third form of legal activity identified by Hildebrandt is “code-driven law”, for which provisions from various legal sources (e.g. legal norms, regulations, policies, agreements, contracts) have been operationalized via computational means. From a historical point of view, the idea of using machines with normative provisions traditionally dates back to Leibniz (Calculemus!), and took more concrete forms with the advent of legal expert systems and various attempts at formalisation of law (e.g. [12, 13]), in part renewed during the semantic web technology heyday, and, in narrower forms, in contemporary RegTech (e.g. smart contracts), and cyber infrastructure research (e.g. data governance, consent management systems). Most distinctive assumptions of these contributions can be categorized within the domain of legal isomorphism [14], susceptible to concerns on whether, to what extent, and how legal sources can be represented in computational artifacts, and normative reasoning automated by computational processes. At a strictly theoretical level, neglecting interpretative and tractability challenges, the problem becomes a purely logical one. Given a formal language, and the associated derivation mechanisms, soundness and completeness can be formally proven to verify a logical closure guaranteeing valid inferences. Let us then suppose that our system is provided with structural knowledge relative to norms in the form of rules, and contingent knowledge relative to the case in the form of facts; within a formal system we can be certain that we utilize these resources to derive logically valid conclusions. But are they correct in a practical sense? Theory and practice of computational legal theory acknowledge several critical questions in this respect: Were all the relevant fact and rules considered? Is the formalization correct? Is the reasoning tool robust? As in the previous section, only the last question concerns proper technical dimensions (e.g. tractability, verification, security); the other two concern the selection of relevant contextual elements, and the choice of mapping from natural to formal language.
@@ 2.4 Setting an alternative technological agenda
Taking into account the critical questions identified in the respective fields of practice, we can conclude that neither data-driven nor code-driven law guarantee legal certainty. The premises on which the underlying computational devices are constructed and deployed will always be “weak” to some extent. Besides, software bugs, limitations or hacks (malicious attacks) may occur in computational systems supporting both data-driven and code-driven law. Said differently, the idea of computational legalism [^4] We refer to Diver [3] for a detailed inquiry on the concept of computational legalism. Related concepts can also be found in early works in AI & Law, cf. mechanical jurisprudence in [15], or the “legalistic child” example discussed in [16]. works only in so far we forget all what may go wrong. The only way by which these computational devices may achieve legal certainty is by removing the possibility of contestation, which, although at times this is presented as a feature, it clearly brings risks of societally detrimental effects.
@@
Perhaps, a more sustainable path is to reproduce in the computational realm procedures analogous to the ones introduced in human societies, which have proven value to deal with the legal (un)certainty of text-guided law. For instance, in order to satisfy principles of justice, such as proportionality and distribution, we may formalize those as meta-rules. In order to keep instrumentality with respect to contextual policy goals, we may require to make these goals explicit, and have them in the loop of the automated reasoning process. The resulting (naive) computational justice stance is based on the belief that a perfect conceptualization will solve all the observed issues. Unfortunately, there are two serious hindrances for such an approach. First, we cannot be sure of any formalization, ever. The strongest support for this claim is not technical, but comes from the human dimension of the problem, the “referent” (the entity referred to) that needs to be formalized. For instance, the debates on ethical frameworks and human rights make clear that there is no such a referent, or, if it exists, it does so only on some contextual basis. [^5] See e.g. Schlag’s metaphor of law as “hidden ball” [17]. Second, we may never be ready to have an explicit and complete articulation of policy-makers’ purposes. Even imaging that this was possible, by means of perfect introspection and no errors in expression, making clear in detail all what some authority aims to achieve (or does not aim to) would entail to be transparent also on highly contestable issues, which in many occasions would be politically detrimental, and thus not to be expected.
@@
When both computational legalism and (naive) computational justice have been defeated, it seems there is no much hope for an advance in legal technologies. Fortunately, this is not true; new and sounder objectives can be set by starting from opposite assumptions. With respect to computational legalism, we should not suppose that things are correct by default (or eventually will be), but instead accept that things may always go wrong, at all levels. Under this new expectation, we may still accept to use legal technologies, as long as we also accept that people or other actors ought to be able to appeal algorithmic decisions. Similarly, with respect to computational justice, we should accept that there may be always relevant directives not adequately articulated in the present computational artifacts. This entails that we need systems to integrate systematically feedbacks from higher-order courts, from jurisprudence or other competent normative sources. In order words, courts or other actors ought to be able to quash or overrule algorithmic decision-making processes. Accepting these two principles leads us to technological innovations which have been neglected so far, aiming to address primarily computational instruments that operationalize appeal and quashing/overruling within the computational realm. Interestingly, such an infrastructure would support the systematical introduction of artificial agents, additionally to humans, to perform critical tasks like testing and auditing.
In short, the core challenge of regulatory technologies is not captured only by requirements as those expressed in software engineering terms, but by the construction of proper, functioning feedback loops, in the form of intervention points for social actors provided with relevant competence. This level of abstraction makes clear that we are in fact in front of a problem of governance (who has the power to do what, under which circumstances, and for what purpose), a problem which, when computational systems are involved as mediators, has also a technical dimension to be dealt with. This is where the concept of normware fits into.